OpenStudy (anonymous):

Simplify the following radical: (6+4i)(1+3i)

OpenStudy (agreene):

start by expanding the factors, and then simplify like terms. keep in mind the rule \(i^2=-1\)

OpenStudy (anonymous):

do FOIL then combine like terms in a+bi

OpenStudy (anonymous):

here is my two cents you can "foil" (ugh) if you like, but the way to multiply two complex numbers is \[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\]

OpenStudy (anonymous):

So, 6+18i+(-12)+(-12)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

6×1 6×3i 4i×1 4i×3i

OpenStudy (anonymous):

\[i ^{2}=-1\]

OpenStudy (anonymous):

Call me stubborn, but is it something like 6-(-12)+(-22)?

OpenStudy (anonymous):

you got the 6 right

OpenStudy (anonymous):

Alright, here is what I have written down.

OpenStudy (anonymous):

(6*1-12i)+(18i+4i)i

OpenStudy (anonymous):

I have a feeling that when I multiplied I shouldn't have switched the 12 to a negative.

OpenStudy (anonymous):

Nor the others.

OpenStudy (anonymous):

Do 6×1 6×3i 4i×1 4i×3i

OpenStudy (anonymous):

after that ill help combine like terms

OpenStudy (anonymous):

I did, and for those I got the following. 6 18i 4i -12

OpenStudy (anonymous):

okay now combine like terms in A+bi form 6-12 18i+4i

OpenStudy (anonymous):

-6+22i?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

That is as low as it can go I assume?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Alright I have a few more questions if you wouldn't mind sticking around.

OpenStudy (anonymous):

I can try to help

OpenStudy (anonymous):

(2-5i)(7+i)?

OpenStudy (anonymous):

use foil for this one too 2×7 2×i -5i×7 -5i×i

OpenStudy (anonymous):

14 2i -35i 5?

OpenStudy (anonymous):

yes now combing like terms in a+bi form 14+5 2+(-35i)

OpenStudy (anonymous):

19-33i?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you! Now, one last one. (-3-6i)(4-2i)

OpenStudy (anonymous):

-3×4 -3×-2i -6i×4 -6i×-2i remember 2 negatives will make a positive

OpenStudy (anonymous):

So: -12 6i -24i 12?

OpenStudy (anonymous):

wouldnt be 12 remember multiplying those will make it positive the i^2=-1 so you multiply negative by positive

OpenStudy (anonymous):

So -12?

OpenStudy (anonymous):

yes -12-12 6i-24i

OpenStudy (anonymous):

Danke sehr!

OpenStudy (anonymous):

And one factoring question: 5x^2-30x=0

OpenStudy (anonymous):

ewww factoring lol

OpenStudy (anonymous):

umm hold on ill look at it

OpenStudy (anonymous):

Heh.

OpenStudy (anonymous):

This...will prove VERY USEFUL.

OpenStudy (anonymous):

Danke und auf wiedersehen!